ADNAN MUHAMMED

Grade 12,

In a rocket, the mass of the fuel is 90% of the total mass.The exhaust gases are ejected at a speed of 2 km/s. Find the maximum speed attained by the rocket. Neglect gravity and air resistance and take the initial speed of the rocket to be zero.

In a rocket, the mass of the fuel is 90% of the total mass.The exhaust gases are ejected at a speed of 2 km/s. Find the maximum speed attained by the rocket. Neglect gravity and air resistance and take the initial speed of the rocket to be zero.

Grade:11

1 Answers

Vikas TU
7 years ago
F = dP/dt = > d(mv)/dt = > vdm/dt + mdv/dt
after integrating something we get,
v = uln(mi/m) + vi
 = 2000 * 2.303*log(0.9)
 = – 210.75 m/s
in opposirte direction of that fuel eecting.